### v1

@@ 5 Federated Learning’s Limitations and Difficulties
Federated Learning (FL) has several advantages, but its complete adoption across industries is hampered by a number of obstacles, particularly those related to privacy, security, and technical constraints. The fact that FL training data is inherently imperfect—it might be biased, uneven, or incomplete—is one of the main obstacles [49]. Poor model performance results from an uneven distribution of training samples among entities, a phenomenon known as data imbalance [50]. The training process is made more difficult by missing classes, features, and values since distinct entities may have datasets that are missing crucial information, which leads to erroneous models. Moreover, the complexity is increased by the heterogeneity of data resulting from its dispersion across several places, rendering crude applications of FL models ineffectual.
@@ 6 Conclusion
Federated Learning (FL), which decentralizes the training process across multiple clients while preserving data security and privacy, offers a revolutionary approach to machine learning. This article explored the fundamental ideas of FL in great detail, assessed the efficacy of several federated learning algorithms, and looked closely at the numerous architectures that make up the FL ecosystem. The substantial advantages of FL, including improved privacy, regulatory compliance, and cost savings, are highlighted by our analysis. But FL also has to deal with a number of issues that call for creative solutions, like communication overhead, data heterogeneity, and privacy problems. The objective of this work is to make a contribution to the progress of FL research and its application in many fields by tackling these restrictions and investigating possible future paths. We also put forward a set of architectural principles that can direct the creation and execution of reliable and effective federated learning systems through a thorough analysis of the literature. This work opens the door for more investigation and advancement in this exciting topic by offering a useful resource for comprehending the fundamental elements and architectures of FL.

### v2

@@ 6 Federated Learning’s Limitations and Difficulties
FL has several advantages, but its complete adoption across industries is hampered by a number of obstacles, particularly those related to privacy, security, and technical constraints. The fact that FL training data is inherently imperfect—it might be biased, uneven, or incomplete—is one of the main obstacles [51]. Poor model performance results from an uneven distribution of training samples among entities, a phenomenon known as data imbalance [52]. The training process is made more difficult by missing classes, features, and values since distinct entities may have datasets that are missing crucial information, which leads to erroneous models. Moreover, the complexity is increased by the heterogeneity of data resulting from its dispersion across several places, rendering crude applications of FL models ineffectual.
@@
7 Future Research on FL and Zero Trust Security
Combining Federated Learning (FL) with Zero Trust Architecture (ZTA) can greatly improve security, privacy, and efficiency in distributed AI systems. To ensure that no one party controls the training process completely, future research should look into using blockchain technology in FL. This would create secure, decentralized records for model updates, reducing the risk of harmful model changes and increasing trust and transparency.
Furthermore, adding differential privacy to FL-based ZTA can help protect sensitive information when model updates are shared. This approach is especially important in fields like healthcare, finance, and government, where data privacy and legal compliance are critical.
Additionally, real-time adaptive authentication using FL can improve security and access control by continuously checking users and devices based on their behavior patterns. By exploring these areas, FL-enabled ZTA can help create a safer, more resilient, and privacy-focused AI environment.
@@ 8 Conclusion
FL transforms machine learning by decentralizing training among clients while ensuring data security and privacy. This article examines the core concepts of FL, evaluates various federated learning algorithms, and analyzes the architectures within the FL ecosystem. Our findings highlight FL’s significant benefits, including enhanced privacy, regulatory compliance, and cost reductions, while also addressing challenges such as communication overhead, data heterogeneity, and privacy issues. The goal of this work is to advance FL research and applications by tackling these challenges and exploring future directions. We also present architectural principles to guide the development of effective federated learning systems, providing a resource for understanding the fundamental aspects of FL and paving the way for further advancements in this area.